OpenStudy (anonymous):

Medal and Fan!!!! What value of s makes the statement true? (-5 + s) + (-15) = (8 + (-5)) + (-15) A. -8 B. 8 C. -15 D. -5

syed98 (syedmohammed98):

its B. :)

OpenStudy (anonymous):

help with another?

syed98 (syedmohammed98):

another?

OpenStudy (anonymous):

What is the value of the expression? -(9 + (-12.9)) + (-3.6) A. 18.3 B. -7.5 C. 0.3 D. 12.3

syed98 (syedmohammed98):

its C :)

OpenStudy (anonymous):

i got 4 more

syed98 (syedmohammed98):

no problem my friend :)

OpenStudy (anonymous):

What value of w makes the statement true? w + (-8.2) = -5.9 A. 2.3 B. 14.1 C. -2.3 D. 12.6

syed98 (syedmohammed98):

-5.9 + 8.2 = 2.3 its A :)

OpenStudy (anonymous):

A

OpenStudy (anonymous):

What value of b makes the statement true? 18 = b - (-3.5) A. 21.5 B. -20.5 C. 5.5 D. 14.5

syed98 (syedmohammed98):

18 - 3.5 = 14.5 its D

OpenStudy (anonymous):

On a number line, point P has coordinate -1.2 and point Q has coordinate -5.8. Which expressions represent the distance between point P and point Q? Choose all answers that are correct. A. |-1.2 + (-5.8)| B. |-1.2 - (-5.8)| C. |-5.8 - (-1.2)| D. |-5.8 - 1.2|

syed98 (syedmohammed98):

i think its C -5.8 - (-1.2)

OpenStudy (anonymous):

The temperature at noon was -6.9°C. The temperature rose 2.8°C from noon to 1 p.m. The temperature increased t degrees centigrade from 1 p.m. to 2 p.m. The temperature at 2 p.m. was -3°C. Which equation can be used to solve for t? A. -6.9 + 2.8 + t = -3 B. -6.9 -2.8 + t = -3 C. -6.9 -2.8 - t =-3 D. -6.9 + 2.8 - t = -3

syed98 (syedmohammed98):

its A

OpenStudy (anonymous):

What is the value of -2.3 - q when q = 7.2? A. 9.5 B. -9.5 C. 5.1 D. -4.9

syed98 (syedmohammed98):

its -2.3 - (7.2) = -9.5 its B

OpenStudy (anonymous):

thanks man wanna help me with another test